If height is 50meter what is the horse supplied by the driving motor to pump tack 1L ~1kg calculate Hp

Grade:12th pass

1 Answers

Khimraj
6 years ago
Work done by motor per sec W = mgh = 1*10*50 = 500J
1HP = 746J/s
then HP supplied to driving motor = 500/746 = 0.67HP
Hope it clears.